OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)? Q(1, 11) R(2, 4) S(4, -4) T(9, -2)

OpenStudy (anonymous):

R

OpenStudy (anonymous):

The equation of a circle is: \[(x-a)^2+(y-b)^2=r^2\] Here, a=6, b=1 and r=5 From this, you can work out the equation of the circle. Then you can put in the different (x,y) values for the choices you have and see which one gives the correct \[r^2\]